### v1

# Traffic Scenario Logic: A Spatial-Temporal Logic for Modeling and Reasoning of Urban Traffic Scenarios

Abstract
Formal representations of traffic scenarios can be used to generate test cases for the safety verification of autonomous driving. However, most existing methods are limited in highway or highly simplified intersection scenarios due to the intricacy and diversity of traffic scenarios. In response, we propose Traffic Scenario Logic (TSL), which is a spatial-temporal logic designed for modeling and reasoning of urban pedestrian-free traffic scenarios. TSL provides a formal representation of the urban road network that can be derived from OpenDRIVE, i.e., the de facto industry standard of high-definition maps for autonomous driving, enabling the representation of a broad range of traffic scenarios. We implemented the reasoning of TSL using Telingo, i.e., a solver for temporal programs based on the Answer Set Programming, and tested it on different urban road layouts. Demonstrations show the effectiveness of TSL in test scenario generation and its potential value in areas like decision-making and control verification of autonomous driving.
1 Introduction
Autonomous driving technology is promising to revolutionize transportation by enhancing safety, efficiency, and accessibility. The rigorous testing and verifying their safety under various traffic scenarios is central to developing and deploying autonomous vehicles. Formal representations of these scenarios are crucial for generating comprehensive test cases and ensuring the reliability of autonomous driving systems. (Zhang et al., 2021)
Existing methods for formalizing traffic scenarios often face limitations, particularly in modeling intricate urban environments with diverse traffic patterns. While some approaches excel in representing highway scenarios or simplified intersections, they struggle to capture the complexity inherent in urban driving environments. (Hilscher and Schwammberger, 2016; Yao and Li, 2023; Li, Tao, and Wotawa, 2020) Consequently, there is a pressing need for a versatile and robust formal logic framework capable of modeling a broad spectrum of urban traffic scenarios.
In response to this challenge, we introduce Traffic Scenario Logic (TSL), a novel spatial-temporal logic specifically tailored for modeling and reasoning about urban pedestrian-free traffic scenarios. Unlike previous methods, TSL is designed to provide a comprehensive formal representation of the urban road network, enabling the modeling of various traffic configurations encountered in complex urban environments.
One of the main benefits of TSL is its compatibility with OpenDRIVE, which is the industry-standard format used for high-definition (HD) maps in autonomous driving. (Dupuis, Strobl, and Grezlikowski, 2010) By utilizing OpenDRIVE data, TSL seamlessly translates real-world road networks into logical representations, enabling the representation of a wide range of urban traffic scenarios.
TSL also offers a spatial representation that sets it apart from previous methods. Unlike other methods that use a simple discretized position description, TSL represents the longitudinal position of a vehicle by its positional relationship with other cars. This is inspired by homotopy classes in trajectory planning, where a homotopy class is the set of all possible trajectories from a start to a goal point that avoids obstacles in the same manner. Similarly, we can categorize two scenarios into the same homotopy class if the positional relationships of vehicles in them vary in the same way. Distinguishing between scenarios of different homotopy classes ensures comprehensiveness, while merging same-class scenarios improves efficiency.
We first give the syntax and semantics of TSL for uni-direction multi-lane highway scenarios in Section 4, then extend it to urban scenarios in Section 5. Finally, we implement the reasoning of TSL with Telingo (Cabalar et al., 2019), a solver for temporal programs based on the Answer Set Programming. Section 6 demonstrates several examples on different urban road layouts, showing the effectiveness of TSL in test scenario generation and potential applications in decision-making and control verification of autonomous driving.
In summary, this paper makes the following contributions:
  • Introduces Traffic Scenario Logic (TSL) for formal modeling and reasoning of urban pedestrian-free traffic scenarios.
  • Provides a formal representation of an urban road network that is compatible with the OpenDRIVE HD map format and a spatial representation that corresponds to homotopy classes of scenarios.
  • Implements the reasoning mechanism with Telingo solver and provides application examples on different road layouts.
2 Related Work
Constructing a generic spatial-temporal model for the physical world is a challenging and significant field of study. Prior research has primarily focused on developing different combinations of temporal and spatial logics. Kontchakov et al. (2007) have provided a comprehensive summary of this research.
Hilscher et al. (2011) were one of the first few work on using spatial-temporal logic to model traffic scenarios, proposing Multi-Lane Spatial Logic (MLSL) which provides an abstract model of a uni-directional multi-lane motorway. In MLSL, vehicles’ position, speed, and acceleration constitute a traffic snapshot, and the traffic snapshots at different moments constitute a traffic scenario. Changes between traffic snapshots are described by transition rules, including acceleration and lane changes made by the vehicle, as well as changes in speed and position over time, inspired by $\mathcal{ITL}$ (Moszkowski, 1982). Hilscher, Linker, and Olderog (2013) proposed an extended version of MLSL that supports a bi-directional motorway and length measurement for road segments. Hilscher and Schwammberger (2016) further proposed Urban Multi-Lane Spatial Logic (UMLSL) which contains the modeling of intersections. UMLSL simplifies an intersection of two two-way, single-lane highways into four crossing road segments. Meanwhile, Xu and Li (2016) introduced vertical lanes that intersect with the horizontal lanes by extending the position to two dimensions to create an urban road network, formalized in a grid way. Based on UMLSL, Schwammberger and Alves (2021) proposed Urban Spatial Logic for Traffic Rules (USL-TR) to add static objects and non-autonomous road users. Xu et al. (2019) represented an urban road network with road segments and the arcs that link them together, which was further formalized into a hierarchical structure by Yao and Li (2023). This structure can express complex road networks like roundabouts. Some other research includes modeling an intersection controlled by traffic lights (Loos and Platzer, 2011); using the State-Clock Logic for the representation of time (Bischopink and Olderog, 2022); and employing ontology (Bagschik, Menzel, and Maurer, 2018; Li, Tao, and Wotawa, 2020).
The above methods for describing spatial relationships rely on either road segments or a grid-based occupancy approach, which have their limitations. In most cases, distinguishing between consecutive multiple free grids or road segments is not necessary. On the other hand, there are instances when it is crucial to describe the positional relationships of vehicles within the same grid or road segment. In general, a simple discretized description of the road network is insufficient to meet the needs of test scenario generation.
3 Preliminary
In this section, we provide a brief introduction to the road coordinate system and OpenDRIVE map format, the de facto industry standard of HD maps for autonomous driving.
The road coordinate system, also known as the Frenet frame, describes motions on a continuous differentiable curve in 3-dimensional Euclidean space. Since the slope angle and superelevation of a road in the vertical direction are commonly small, what is used for road coordinate description is mostly a 2-dimensional version. The road coordinate system is composed of the $S$ axis along the direction of the road reference line (generally the road centerline) and its orthogonal $D$ axis (also referred to as the $L$ or $T$ axis in some literature). The coordinates of a point on the road are determined by its accumulated distance $s$ along the road reference line from the start point and the distance $d$ between the point and its projection onto the reference line. With the road coordinate system, we can mitigate the influence of road curvature and uniformly describe the vehicle’s position on the road.
reference line $d$$s$$S$$D$projection point
Figure 1: The road coordinate system. The coordinate of the red dot consists of its accumulated distance $s$ along the reference line and the distance $d$ to the projection point.
OpenDRIVE [^1] https://www.asam.net/standards/detail/opendrive/ is a standardized HD map file format for representing road networks, encapsulating information about road geometry, topology, traffic signs, etc. Each road in an OpenDRIVE file consists of road segments with a specific number of lanes, as depicted in 2(a). Junctions are used to connect the roads, creating a connecting road for each pair of connected lanes. Due to its powerful representation capabilities, OpenDRIVE has become the most widely used HD map format in autonomous driving and simulators such as CARLA.
Junction
(a) OpenDRIVE map
connect pointintersect pointl1l2l3l4l5l6l13l42l45l62l63l15n1n2n3n4n5n6x1x2x3
(b) Topology abstraction.
Figure 2: A typical T-intersection. Six connecting roads link three 2-lane roads in the junction area.
4 Traffic Scenario Logic for Highway
4.1 Abstract Model
This section formally defines a scenario on a uni-direction multi-lane highway, where one or more vehicles travel on a road with multiple lanes. The road length is infinite, and the number of lanes remains constant throughout. Vehicles can perform lane-changing maneuvers between lanes to switch to adjacent lanes.
We denote the set of vehicles as =$\mathbb{C}=\{c_{1},c_{2},\dots,c_{N}\}$, where $N$ is the number of vehicles in the scenario. The set of lanes is denoted as $\mathbb{L}=\{l_{1},l_{2},\dots,l_{M}\}$, where $M$ is the number of lanes. Then, we can define a road as follows: {definition}[Road] A road is an ordered arrangement of lanes of the lane set $\mathbb{L}$, denoted as $\mathcal{R}=\langle l_{i_{1}},l_{i_{2}},\dots,l_{i_{M}}\rangle$. A binary relation $\prec_{\text{left}}$ is defined on it, where $l_{i_{j}}\prec_{\text{left}}l_{i_{j+1}},1\leq j\leq M-1$.
The lateral position of a vehicle (perpendicular to the direction of the road) is described using the lanes it occupies, while the longitudinal position (along the direction of the road) can be described using the positional relationship with other vehicles. Using the road coordinate system, it can be defined as follows: {definition}[Longitudinal Positional Relationship] Denote the range occupied by vehicle $c$ as $[s_{c}^{-},s_{c}^{+}]$, where $s_{c}^{-}$ and $s_{c}^{+}$ represent the $s$-coordinates of the rear and front ends of the vehicle, respectively. Assuming that the $S$ axis direction aligns with the direction of the vehicle, based on the road coordinate system, we can categorize the longitudinal positional relationship between two vehicles $c_{1}$ and $c_{2}$ on the same road as follows:
  1. 1.
    Ahead: If $s_{c_{1}}^{-}>s_{c_{2}}^{+}$, then $c_{1}$ is ahead $c_{2}$.
  2. 2.
    Cover: If $s_{c_{1}}^{-}\leq s_{c_{2}}^{+}$ and $s_{c_{1}}^{+}\geq s_{c_{2}}^{-}$, then $c_{1}$ is covering $c_{2}$.
  3. 3.
    Behind: If $s_{c_{1}}^{+}<s_{c_{2}}^{-}$, then $c_{1}$ is behind $c_{2}$.
The definition is similar when the $S$ axis direction is opposite the direction of the vehicle.
Then we can define a scene, which is a “frame” of a scenario: {definition}[Scene on Highway] A scene on a uni-direction multi-lane highway is a pair $\mathfrak{s}=(\mathscr{L},\mathscr{D})$, where:
    $\mathscr{L}:\mathbb{C}\mapsto\mathcal{P}(\mathbb{L})$ is a mapping from the set of vehicles $\mathbb{C}$ to the power set of the set of lanes $\mathcal{P}(\mathbb{L})$, assigning to each vehicle the set of lanes to which it belongs; note that vehicles may cross lane dividers, so a vehicle may belong to more than one lane.
    $\mathscr{D}:\mathbb{C}\times\mathbb{C}\mapsto\mathbb{D}$ is a mapping from the set of vehicle pairs to the set of longitudinal positional relationship $\mathbb{D}=\{\mathsf{ahead},\mathsf{cover},\mathsf{behind}\}$.
A scenario is a sequence of scenes, defined as follows: {definition}[Scenario on Highway] A scenario on a uni-direction multi-lane highway is a quintuple $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},T,\langle\mathfrak{s}_{0},\mathfrak{s}_{1},\dots,\mathfrak{s}_{T-1}\rangle)$, where:
    The meaning of $\mathbb{C}$, $\mathbb{L}$ and $\mathbb{D}$ is as stated above;
    $T\geq 1$ is the number of scenes included in the scenario, i.e. the length of the scenario;
    $\mathfrak{s}_{i}=(\mathscr{L}_{i},\mathscr{D}_{i})$ is the $i$-th scene that occurs in the scenario in time sequence.
For convenience, further define the tail of a scenario as: {definition}[Tail of a Scenario] The tail of scenario $\mathcal{S}$, denoted by $\mathcal{S}_{i:}=(\mathbb{C},\mathbb{L},\mathbb{D},T-i,\langle\mathfrak{s}_{i},\mathfrak{s}_{i+1},\dots,\mathfrak{s}_{T-1}\rangle)$ $(0\leq i<T)$
Now we provide the syntax and semantics of the TSL for uni-direction multi-lane highway scenario.
{definition}
[Syntax of Highway TSL] A formula of highway TSL can be recursively defined as:
$\begin{gathered}\phi::=\top\mid\mathsf{on}(c,l)\mid\mathsf{left}(l_{1},l_{2})\mid\mathsf{lonr}(c_{1},c_{2},d)\mid c_{1}=c_{2}\\ \mid l_{1}=l_{2}\mid d_{1}=d_{2}\mid\lnot\phi\mid\phi_{1}\to\phi_{2}\mid\mdwhtcircle\phi\mid\mdwhtsquare\phi\\ \mid\mathbfbb{F}\mid\forall c\,\phi\mid\forall l\,\phi\mid\forall d\,\phi\end{gathered}$
where, $c,c_{1},c_{2}\in\mathbb{C}$, $l,l_{1},l_{2}\in\mathbb{L}$, $d,d_{1},d_{2}\in\mathbb{D}$. The modal operators and take their meanings “Next” and “Globally” in temporal logic, respectively. $\mathbfbb{F}$ denotes that the length of this scenario is 1, indicating the final state. {remark} We can further define $\phi\land\psi$ , $\phi\lor\psi$ , $\exists x\,\phi$ and $\mdwhtlozenge\phi$ as ($(\lnot(\phi\to(\lnot\psi)))$ , ((¬$((\lnot\phi)\to\psi)$ , ($(\lnot\forall x\,(\lnot\phi))$ and ($(\lnot\mdwhtsquare\lnot\phi)$.
Please note that the temporal modal operators here represent the transition of scenes, indicating only the chronological order, not the exact time. The interval between steps can last for an arbitrary duration and can be different.
{definition}
[Semantics of Highway TSL] Given a road $\mathcal{R}$ and a scenario $\mathcal{S}$, the satisfaction of formulae in the subsection 4.1 follows the semantics listed below:
[…]
$\mathcal{R},\,\mathcal{S}\models\mathsf{on}(c,l)$$\Leftrightarrow$ $l\in\mathscr{L}_{0}(c)$
$\mathcal{R},\,\mathcal{S}\models\mathsf{left}(l_{1},l_{2})$$\Leftrightarrow$ $l_{1}\prec_{\text{left}}l_{2}$
$\mathcal{R},\,\mathcal{S}\models\mathsf{lonr}(c_{1},c_{2},d)$$\Leftrightarrow$ $\mathscr{D}_{0}(c_{1},c_{2})=d$
$\mathcal{R},\,\mathcal{S}\models\lnot\phi$$\Leftrightarrow$ $\mathcal{R},\,\mathcal{S}\not\models\phi$
$\mathcal{R},\,\mathcal{S}\models\phi_{1}\to\phi_{2}$$\Leftrightarrow$ $\mathcal{R},\,\mathcal{S}\models\phi_{2}\;\text{if}\;\mathcal{R},\,\mathcal{S}\models\phi_{1}$
$\mathcal{R},\,\mathcal{S}\models\mdwhtcircle\phi$$\Leftrightarrow$ $T>1\;\text{and}\;\mathcal{R},\,\mathcal{S}_{1:}\models\phi$
$\mathcal{R},\,\mathcal{S}\models\mdwhtsquare\phi$$\Leftrightarrow$ $\forall 0\leq i<T,\,\mathcal{R},\,\mathcal{S}_{i:}\models\phi$
$\mathcal{R},\,\mathcal{S}\models\mathbfbb{F}$$\Leftrightarrow$ $T=1$
4.2 Rules of Highway TSL
Realistic scenarios follow physics rules and traffic rules. In this subsection, we list these rules in the form of TSL formulae, where the traffic rules can be added or subtracted according to the laws of different regions.
Physics Rules
  1. 1.
    The longitudinal positional relationships are symmetric
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mathsf{lonr}(c_{2},c_{1},\mathsf{behind})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\to\mathsf{lonr}(c_{2},c_{1},\mathsf{ahead})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\to\mathsf{lonr}(c_{2},c_{1},\mathsf{cover})).$
  2. 2.
    The longitudinal positional relationships are transitive.
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{ahead})\to\mathsf{lonr}(c_{1},c_{3},\mathsf{ahead})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{behind})\to\mathsf{lonr}(c_{1},c_{3},\mathsf{behind})).$
  3. 3.
    Contradictory longitudinal positional relationships.
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\lnot(\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{cover})\land\mathsf{lonr}(c_{3},c_{1},\mathsf{ahead}))),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall c_{3}\,\lnot(\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\\ \land\mathsf{lonr}(c_{2},c_{3},\mathsf{ahead})\land\mathsf{lonr}(c_{3},c_{1},\mathsf{ahead}))).$
  4. 4.
    The longitudinal positional relationships must change continuously.
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\\ \to\mdwhtcircle(\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\lor\mathsf{lonr}(c_{1},c_{2},\mathsf{cover}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c_{1}\forall c_{2}\,\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\\ \to\mdwhtcircle(\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\lor\mathsf{lonr}(c_{1},c_{2},\mathsf{cover}))).$
  5. 5.
    The occupied lanes must be continuous, meaning that the lanes occupied by a vehicle must be one or more adjacent lanes. The logical representation of this rule is slightly complex. First, we introduce an auxiliary predicate $\mathsf{cleft/2}$, representing one lane being to the left of another lane but not necessarily adjacent. This can be achieved by defining it recursively as follows:
    $\mdwhtsquare(\forall l_{1}\forall l_{2}\,\mathsf{left}(l_{1},l_{2})\to\mathsf{cleft}(l_{1},l_{2})),$
    $\mdwhtsquare(\forall l_{1}\forall l_{2}\forall l_{3}\,\mathsf{left}(l_{2},l_{3})\land\mathsf{cleft}(l_{1},l_{2})\to\mathsf{cleft}(l_{1},l_{3})).$
    Therefore, the rule stating that the occupied lanes must be continuous can be expressed as:
    $\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\forall l_{3}\,\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\\ \land\mathsf{cleft}(l_{1},l_{3})\land\mathsf{cleft}(l_{3},l_{2})\to\mathsf{on}(c,l_{3})),$
    which means any lane between two lanes occupied by a vehicle must also be occupied by that vehicle.
  6. 6.
    Each vehicle always occupies at least one lane.
    $\mdwhtsquare(\forall c\forall l\,\exists l\mathsf{on}(c,l)).$
  7. 7.
    The occupied lanes must change continuously. That is, only one lane can be added to or removed from the set of occupied lanes on each step.
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\\ \land\lnot(l_{1}=l_{2})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{1})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{2}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\lnot\mathsf{on}(c,l_{1})\land\lnot\mathsf{on}(c,l_{2})\\ \land\lnot(l_{1}=l_{2})\land\mdwhtcircle\mathsf{on}(c,l_{1})\land\mdwhtcircle\mathsf{on}(c,l_{2}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall l_{1}\forall l_{2}\,\lnot(\mathsf{on}(c,l_{1})\land\lnot\mathsf{on}(c,l_{2})\\ \land\lnot(l_{1}=l_{2})\land\lnot\mdwhtcircle\mathsf{on}(c,l_{1})\land\mdwhtcircle\mathsf{on}(c,l_{2}))).$
Traffic Rules
  1. 1.
    Vehicles are not allowed to occupy three or more lanes simultaneously.
    $\mdwhtsquare\forall c\forall l_{1}\forall l_{2}\forall l_{3}\,\lnot(\mathsf{on}(c,l_{1})\land\mathsf{on}(c,l_{2})\land\mathsf{on}(c,l_{3})\\ \land\lnot(l_{1}=l_{2})\land\lnot(l_{1}=l_{3})\land\lnot(l_{2}=l_{3})).$
  2. 2.
    Two vehicles cannot drive side by side in the same lane.
    $\mdwhtsquare\forall c_{1}\forall c_{2}\forall l\,\lnot(\mathsf{on}(c_{1},l)\land\mathsf{on}(c_{2},l)\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\land\lnot(c_{1}=c_{2})).$
A realistic scenario must satisfy all the aforementioned rules. Consequently, by utilizing these rules, we can validate the plausibility of a scenario, rendering the reasoning of scenarios under TSL executable.
5 Traffic Scenario Logic for Urban
In this section, we will introduce the extension of TSL to general urban traffic scenarios, including vehicles traveling on structured roads without pedestrians.
5.1 Formal Representation of Urban Road Network
A structured urban road network can be described by an OpenDRIVE map, as shown in Figure 2. Considering two lanes on the ground, their relationships can be categorized as follows:
  1. 1.
    They belong to the same road, in which case the scenario is identical to that described in Section 4;
  2. 2.
    They have an intersection point, but are not connected at this point;
  3. 3.
    They are connected at a connection point, where vehicles can switch to another lane;
  4. 4.
    They have an overlap segment in which vehicles travel in the same direction. In this case, it can be considered as one lane;
  5. 5.
    They have an overlap segment in which vehicles travel in opposite directions;
  6. 6.
    They are unrelated.
All we need is to extend the TSL for intersections, connections, and overlapping in opposite directions. The rest of the cases can reuse the logical representation defined in Section 4.
  1. 1.
    Intersection
    Without loss of generality, we only consider the situation where two lanes intersect once. For multiple intersections, we can simply apply it repeatedly to each intersection.
    Denote the intersection point as $p^{\mathsf{x}}$. Then, for any vehicle $c$ on these two intersecting lanes, there are essentially three positional relationships with respect to this intersection point: yet to pass, currently passing, and already passed. We use the predicate $\mathsf{lonpr/3}$ to represent these relationships, corresponding to the predicate $\mathsf{lonr/3}$ for vehicles’ relationships. The same predicate can also be used for the connecting points and endpoints of overlapping segments below.
  2. 2.
    Connection
    Two or more lanes can connect at their endpoints. We denote the connecting points as $p^{\mathsf{c}}$. Depending on the direction of travel of vehicles on the lanes, there exists a successor relationship $\succ_{\mathsf{c}}$ between the connecting points and the lanes.
  3. 3.
    Overlapping in opposite directions
    When examining the overlapping segment between points $p^{\mathsf{os}}$ and $p^{\mathsf{oe}}$, as shown in Figure 3, we can designate one of the two overlapping lanes’ direction as the reference direction, i.e. from $p^{\mathsf{os}}$ to $p^{\mathsf{oe}}$. Another longitudinal positional relationship, denoted as $\mathsf{lonro/3}$, exists between the vehicles in the overlapping segment, and:
    • For the lane in the reference direction, the relationships $\mathsf{lonro/3}$ between vehicles are the same as $\mathsf{lonr/3}$;
    • For the lane in the opposite direction, the relationships $\mathsf{lonro/3}$ between vehicles are opposite to $\mathsf{lonr/3}$;
    • For two vehicles belonging to different directions, the relationship $\mathsf{lonro/3}$ between them can be arbitrary, but only monotonic changes are allowed.
$p^{\mathsf{os}}$$p^{\mathsf{oe}}$
Figure 3: A overlapping segment of two lanes with opposite directions
Besides, when there are multiple points (including intersection points, connection points, and endpoints of overlapping segments) on a lane $l$, there also exists a successor relationship $\succ_{\mathsf{p}}^{l}$ between points according to the direction of vehicle travel on the lane.
Based on the above analysis, we further provide a formal definition of the road network:
{definition}
[Road Network] The road network is a nonuple$\mathcal{M}=(\mathbb{R},\mathbb{P}^{\mathsf{X}},\mathbb{P}^{\mathsf{C}},\mathbb{P}^{\mathsf{OS}},\mathbb{P}^{\mathsf{OE}},\mathbb{S}^{\mathsf{P}},\mathbb{S}^{\mathsf{C}},\mathbb{O},\mathbb{T})$, where:
    =$\mathbb{R}=\{\mathcal{R}_{1},\mathcal{R}_{2},\dots\}$ is the set of all roads;
    $\mathbb{P}^{\mathsf{X}},\mathbb{P}^{\mathsf{C}},\mathbb{P}^{\mathsf{OS}},\mathbb{P}^{\mathsf{OE}}$are sets of points $p^{\mathsf{x}}$, $p^{\mathsf{c}}$, $p^{\mathsf{os}}$, and $p^{\mathsf{oe}}$ respectively;
    $\mathbb{S}^{\mathsf{P}}\subseteq\mathbb{L}\times\mathbb{P}\times\mathbb{P}$ defines the successor relationship $\succ_{\mathsf{p}}^{l}$ for points on each lane $l$, where $\mathbb{L}=\bigcup_{i}\mathbb{L}_{i}$ is the set of all lanes, and $\mathbb{P}=\mathbb{P}^{\mathsf{X}}\cup\mathbb{P}^{\mathsf{C}}\cup\mathbb{P}^{\mathsf{OS}}\cup\mathbb{P}^{\mathsf{OE}}$ is the set of all points;
    $\mathbb{S}^{\mathsf{C}}\subseteq\mathbb{P}^{\mathsf{C}}\times\mathbb{L}$ defines the successor lane for connection points, denoted as the relation $\succ_{\mathsf{c}}$;
    $\mathbb{O}\subseteq\mathbb{P}^{\mathsf{OS}}\times\mathbb{P}^{\mathsf{OE}}$ represents the pairs $(p^{\mathsf{os}},p^{\mathsf{oe}})$ for each overlapping segment.
    $\mathbb{T}\subseteq\mathbb{L}\times\mathbb{P}$ defines the affiliation relationship between points and lanes.
2(b) shows the topology abstraction of a typical T-intersection based on Figure 3. There are 12 lanes in it, with 6 connection points (marked by blue diamonds) and 3 intersection points (marked by red circled times).
5.2 Abstract Model
In this subsection, we provide the syntax and semantics of TSL for urban traffic scenarios.
{definition}
[Scene] A (urban) scene is a quadruple $\mathfrak{s}=(\mathscr{L},\mathscr{D},\mathscr{P},\mathscr{Q})$, where:
    $\mathscr{L}:\mathbb{C}\mapsto\mathcal{P}(\mathbb{L})$ is a mapping from the set of vehicles to the power set of the set of lanes, assigning to each vehicle a set of lanes to which it occupies;
    $\mathscr{D}:\mathbb{C}\times\mathbb{C}\mapsto\mathbb{D}$ is a mapping from pairs of vehicles to the set of longitudinal positional relationships $\mathbb{D}=\{\mathsf{ahead},\mathsf{cover},\mathsf{behind},\mathsf{none}\}$, where $\mathsf{none}$ indicates that there is no relationship between two vehicles, which occurs when the two vehicles belong to different roads;
    $\mathscr{P}:\mathbb{C}\times\mathbb{P}\mapsto\mathbb{D}$ is a mapping from pairs of vehicles and points to the set of longitudinal positional relationships, as explained in subsection 5.1;
    $\mathscr{Q}:\mathbb{C}\times\mathbb{C}\mapsto\mathbb{D}$ is a mapping to define the temporary longitudinal positional relationship $\mathsf{lonro/3}$ on the overlapping segments.
{definition}
[Scenario] A (urban) scenario of is a sextuple $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},\mathbb{P},T,\langle\mathfrak{s}_{0},\mathfrak{s}_{1},\dots,\mathfrak{s}_{T-1}\rangle)$, where:
    $\mathbb{C}$, $\mathbb{L}$, $\mathbb{D}$, and $\mathbb{P}$ have the meanings as described above;
    $T\geq 1$ is the length of the scenario;
    $\mathfrak{s}_{i}=(\mathscr{L}_{i},\mathscr{D}_{i},\mathscr{P}_{i},\mathscr{Q}_{i})$ is the $i$-th scene that occurs in the scenario in time sequence.
The definition of the tail of a scenario is similar to subsection 4.1 and will not be repeated.
{definition}
[Syntax of Urban TSL] A formula of urban TSL can be recursively defined as follows:
$\begin{gathered}\phi::=\top\mid\mathsf{on}(c,l)\mid\mathsf{pon}(p,l)\mid\mathsf{left}(l_{1},l_{2})\mid\mathsf{lonr}(c_{1},c_{2},d)\\ \mid\mathsf{lonpr}(c,p,d)\mid\mathsf{lonro}(c_{1},c_{2},d)\mid\mathsf{succp}(l,p_{1},p_{2})\\ \mid\mathsf{succl}(p,l)\mid\mathsf{belong}(l,\mathcal{R})\mid\mathsf{overlap}(p_{1},p_{2})\mid\mathsf{p_{x}}(p)\mid\mathsf{p_{c}}(p)\\ \mid\mathsf{p_{os}}(p)\mid\mathsf{p_{oe}}(p)\mid c_{1}=c_{2}\mid l_{1}=l_{2}\mid d_{1}=d_{2}\mid p_{1}=p_{2}\\ \mid\mathcal{R}_{1}=\mathcal{R}_{2}\mid\lnot\phi\mid\phi_{1}\to\phi_{2}\mid\mdwhtcircle\phi\mid\mdwhtsquare\phi\mid\mathbfbb{F}\\ \mid\forall c\,\phi\mid\forall l\,\phi\mid\forall d\,\phi\mid\forall p\,\phi\mid\forall\mathcal{R}\,\phi\end{gathered}$
{definition}
[The Semantics of Traffic Scenario Logic for Urban] Given the road network $\mathcal{M}$ and the logical scenario $\mathcal{S}$, the satisfaction of formulas defined in subsection 5.2 follows the following semantics:
[…]
$\mathcal{M},\,\mathcal{S}\models\mathsf{on}(c,l)$$\Leftrightarrow$ $l\in\mathscr{L}_{0}(c)$
$\mathcal{M},\,\mathcal{S}\models\mathsf{pon}(p,l)$$\Leftrightarrow$ $(l,p)\in\mathbb{T}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{left}(l_{1},l_{2})$$\Leftrightarrow$ $l_{1}\prec_{\text{left}}l_{2}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{lonr}(c_{1},c_{2},d)$$\Leftrightarrow$ $\mathscr{D}_{0}(c_{1},c_{2})=d$
$\mathcal{M},\,\mathcal{S}\models\mathsf{lonpr}(c,p,d)$$\Leftrightarrow$ $\mathscr{P}_{0}(c,p)=d$
$\mathcal{M},\,\mathcal{S}\models\mathsf{lonro}(c_{1},c_{2},d)$$\Leftrightarrow$ $\mathscr{Q}_{0}(c_{1},c_{2})=d$
$\mathcal{M},\,\mathcal{S}\models\mathsf{succp}(l,p_{1},p_{2})$$\Leftrightarrow$ $(l,p_{1},p_{2})\in\mathbb{S}^{\mathsf{P}}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{succl}(p,l)$$\Leftrightarrow$ $(p,l)\in\mathbb{S}^{\mathsf{C}}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{belong}(l,\mathcal{R})$$\Leftrightarrow$ $l\in\mathcal{R}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{overlap}(p_{1},p_{2})$$\Leftrightarrow$ $(p_{1},p_{2})\in\mathbb{O}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{p_{x}}(p)$$\Leftrightarrow$ $p\in\mathbb{P}^{\mathsf{X}}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{p_{c}}(p)$$\Leftrightarrow$ $p\in\mathbb{P}^{\mathsf{C}}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{p_{os}}(p)$$\Leftrightarrow$ $p\in\mathbb{P}^{\mathsf{OS}}$
$\mathcal{M},\,\mathcal{S}\models\mathsf{p_{oe}}(p)$$\Leftrightarrow$ $p\in\mathbb{P}^{\mathsf{OE}}$
$\mathcal{M},\,\mathcal{S}\models\lnot\phi$$\Leftrightarrow$ $\mathcal{M},\,\mathcal{S}\not\models\phi$
$\mathcal{M},\,\mathcal{S}\models\phi_{1}\to\phi_{2}$$\Leftrightarrow$ $\mathcal{M},\,\mathcal{S}\models\phi_{2}\;\text{if}\;\mathcal{M},\,\mathcal{S}\models\phi_{1}$
$\mathcal{M},\,\mathcal{S}\models\mdwhtcircle\phi$$\Leftrightarrow$ $T>1\;\text{and}\;\mathcal{M},\,\mathcal{S}_{1:}\models\phi$
$\mathcal{M},\,\mathcal{S}\models\mdwhtsquare\phi$$\Leftrightarrow$ $\forall 0\leq i<T,\,\mathcal{M},\,\mathcal{S}_{i:}\models\phi$
$\mathcal{M},\,\mathcal{S}\models\mathbfbb{F}$$\Leftrightarrow$ $T=1$
5.3 Rules of Urban TSL
For urban traffic scenarios, the rules described in subsection 4.2 regarding vehicles within the same road still apply and will not be reiterated here. Additionally, the following rules need to be added to handle situations across lanes:
  1. 1.
    A vehicle can only be located on one road. Firstly, we introduce a auxiliary predicate $\mathsf{cbelong/2}$ to represent the belonging relationship between a vehicle and a road:
    $\mdwhtsquare(\forall c\forall l\forall\mathcal{R}\,\mathsf{belong}(l,\mathcal{R})\land\mathsf{on}(c,l)\to\mathsf{cbelong}(c,\mathcal{R})).$
    Then we have
    $\mdwhtsquare(\forall c\,\forall\mathcal{R}_{1}\forall\mathcal{R}_{2}\,\lnot(\mathsf{cbelong}(c,\mathcal{R}_{1})\land\\ \mathsf{cbelong}(c,\mathcal{R}_{2})\land\lnot(\mathcal{R}_{1}=\mathcal{R}_{2}))).$
  2. 2.
    The longitudinal positional relationship between a vehicle and the points on the same lane can only change monotonically.
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{behind})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{behind})\\ \lor\mathsf{lonpr}(c,p,\mathsf{cover})\lor\mathsf{lonpr}(c,p,\mathsf{none}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{cover})\to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{cover})\\ \lor\mathsf{lonpr}(c,p,\mathsf{ahead})\lor\mathsf{lonpr}(c,p,\mathsf{none}))),$
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\,\mathsf{lonpr}(c,p,\mathsf{ahead})\\ \to\mdwhtcircle(\mathsf{lonpr}(c,p,\mathsf{ahead})\lor\mathsf{lonpr}(c,p,\mathsf{none}))).$
  3. 3.
    A unique longitudinal relationship, which is not $\mathsf{none}$, exists between a vehicle and the points on the lane it occupies.
    $\mdwhtsquare(\forall c\forall p\forall d_{1}\forall d_{2}\,\lnot(\mathsf{lonpr}(c,p,d_{1})\\ \land\mathsf{lonpr}(c,p,d_{2})\land\lnot(d_{1}=d_{2}))),$
    $\mdwhtsquare(\forall c\forall l\forall p\,\lnot(\mathsf{on}(c,l)\land\mathsf{pon}(p,l)\land\mathsf{lonpr}(c,p,\mathsf{none}))).$
  4. 4.
    Only one vehicle is allowed to cover a point.
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p\,\lnot(\mathsf{lonpr}(c_{1},p,\mathsf{cover})\\ \land\mathsf{lonpr}(c_{2},p,\mathsf{cover})\land\lnot(c_{1}=c_{2}))).$
  5. 5.
    When a vehicle passes through a connection point, it will enter one successor lane of that connection point.
    $\mathbfbb{F}\lor\mdwhtsquare(\forall c\forall p\forall l_{1}\,\mathsf{lonpr}(c,p,\mathsf{cover})\land\mathsf{p_{c}}(p)\land\mathsf{on}(c,l_{1})\to\\ \mdwhtcircle((\mathsf{lonpr}(c,p,\mathsf{cover})\land\mathsf{on}(c,l_{1}))\lor\\ \exists l_{2}(\mathsf{lonpr}(c,p,\mathsf{ahead})\land\mathsf{on}(c,l_{2})\land\mathsf{succl}(p,l_{2})))).$
  6. 6.
    Rules for overlapping segments following the discussion in subsection 5.1. Firstly, we define predicates $\mathsf{fwdover/3}$ and $\mathsf{rvsover/3}$ to indicate the vehicle driving along the reference direction and against the reference direction on an overlapping segment, respectively.
    $\mdwhtsquare(\forall c\forall p_{1}\forall p_{2}\forall l\,\mathsf{lonpr}(c,p_{1},\mathsf{ahead})\land\mathsf{lonpr}(c,p_{2},\mathsf{behind})\\ \land\mathsf{on}(c,l)\land\mathsf{pon}(p_{1},l)\land\mathsf{pon}(p_{2},l)\\ \land\mathsf{overlap}(p_{1},p_{2})\to\mathsf{fwdover}(c,p_{1},p_{2})),$
    $\mdwhtsquare(\forall c\forall p_{1}\forall p_{2}\forall l\,\mathsf{lonpr}(c,p_{1},\mathsf{behind})\land\mathsf{lonpr}(c,p_{2},\mathsf{ahead})\\ \land\mathsf{on}(c,l)\land\mathsf{pon}(p_{1},l)\land\mathsf{pon}(p_{2},l)\\ \land\mathsf{overlap}(p_{1},p_{2})\to\mathsf{rvsover}(c,p_{1},p_{2})).$
    Then we have the formulae about $\mathsf{lonro/3}$ as mentioned above:
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\forall d\,\\ \mathsf{fwdover}(c_{1},p_{1},p_{2})\land\mathsf{fwdover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},d)\to\mathsf{lonro}(c_{1},c_{2},d)),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{ahead})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{behind})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{cover})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{cover})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{rvsover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \land\mathsf{lonr}(c_{1},c_{2},\mathsf{behind})\to\mathsf{lonro}(c_{1},c_{2},\mathsf{ahead})),$
    $\mdwhtsquare(\forall c_{1}\forall c_{2}\forall p_{1}\forall p_{2}\,\mathsf{fwdover}(c_{1},p_{1},p_{2})\land\mathsf{rvsover}(c_{2},p_{1},p_{2})\\ \to(\exists d\,\mathsf{lonro}(c_{1},c_{2},d)\land\lnot(d=\mathsf{none}))).$
  7. 7.
    All longitudinal positional relationships, including those between vehicles and vehicles, vehicles and points, and those on overlapping segments, follow similar rules as presented in subsection 4.2, which will not be reiterated here.
By adding the above rules, we can validate the plausibility of an urban traffic scenario.
6 Application
(a) Highway
(b) Intersection
(c) Connecting Roads
(d) A vehicle passes two intersection points
(e) Overtake on the opposite side (overlapping lanes)
Figure 4: Examples of generated scenarios with TSL
TSL is a useful tool for modeling and reasoning traffic scenarios. It can generate various traffic scenarios, which can be used to create a library of test cases for comprehensive testing of autonomous vehicles. (Zhang et al., 2021; Bagschik, Menzel, and Maurer, 2018) Additionally, TSL can identify potential failure risks during the development of autonomous driving strategies, which is our previous work. (Wang et al., 2023) Its ability to differentiate between homotopy classes of scenarios makes it well-suited for these applications, which require both the comprehensiveness and generalization of scenarios.
In this section, we start by introducing the reasoning of TSL, which we implemented using answer set programming with temporal logic extensions. Following that, we illustrate several examples of TSL for scenario generation to demonstrate its effectiveness in handling miscellaneous urban scenarios.
6.1 Reasoning of TSL
Given an initial scene $\mathfrak{s}_{0}$, the set of vehicles $\mathbb{C}$, a finite horizon $T$ and a road network $\mathcal{M}$, the reasoning of TSL is to find all logical scenarios $\mathcal{S}=(\mathbb{C},\mathbb{L},\mathbb{D},\mathbb{P},T,\langle\mathcal{S}_{0},\mathcal{S}_{1},\dots,\mathcal{S}_{T-1}\rangle)$ that $\mathcal{M},\mathcal{S}\models\Phi$, where $\Phi$ is the set of all rules explained in Section 4 and Section 5.
We use Telingo (Cabalar et al., 2019) to implement the reasoning of TSL. Telingo is a solver for temporal logic programs based on the Answer Set Programming (ASP) and the well-known Clingo (Gebser et al., 2019) input language. Generally, a logic program of the Answer Set Programming consists of rules like
A0 :- L1,L2,…,Ln.
where the head A0 is an atom, and the body is the conjunction of literals Li$(1\leq i\leq n)$ of the form A or not A, where A is an atom and not represents the Negation as Failure (NAF). A rule without a body is a fact, and a rule without a head becomes a constraint. Telingo extends the ASP language by importing temporal modal operators under the semantics of the Linear Temporal Logic ($\mathcal{LTL}$(Cabalar et al., 2018; Pnueli, 1977).
We use aggregates (Son and Pontelli, 2007) in the logic program we used for TSL reasoning to derive different scenarios. All rules introduced in Section 4 and Section 5 are rewritten into corresponding ASP rules or constraints. The road network and initial scene are described by facts. For the entire implementation of our reasoning programs, please refer to our source code [^2] https://github.com/Autonomous-Driving-Safety-Project/TSL.
6.2 Examples of Scenario Generation
In this subsection, we provide several examples of TSL on different road layouts.
Example 1 (Highway scenario)
4(a) shows a uni-direction two-lane highway. The road network can be written as:
#program always.
is_road(r1).
is_lane(l1;l2).
has_lane(r1,l1).
has_lane(r1,l2).
left(l1,l2).
The initial scene is:
#program initial.
on_lane(c1,l2).
on_lane(c2,l2).
lonr(c1,c2,behind).
By adding a constraint that c2 shall not be in front of c1 in the final scene:
#program final.
:- lonr(c2,c1,ahead).
We can obtain the 4 shortest overtaking scenarios in 3 steps, where the first one is
State 0:
lonr(c2,c1,ahead) lonr(c1,c2,behind)
on(c1,l2) on(c2,l2)
State 1:
lonr(c2,c1,ahead) lonr(c1,c2,behind)
on(c1,l1) on(c1,l2) on(c2,l2)
State 2:
lonr(c1,c2,cover) lonr(c2,c1,cover)
on(c1,l1) on(c2,l2)
All models we get are shown in 4(a). In the first scenario, c1 (the blue one) overtakes c2 (the red one) by lane change. Two cars change lanes simultaneously in the second and third scenarios, and then one gives up. In the fourth scenario, c2 yields to c1.
Example 2 (Intersection)
When two cars pass an intersection point, TSL gives two different scenarios, as shown in 4(b). Please note that the two lanes are not connected at the intersection point, so a car cannot switch to another lane.
Example 3 (Connecting roads)
4(c) illustrates a scenario where a car passes through a connection point and may drive into any of the connecting lanes.
Example 4 (A vehicle passes two intersection points)
When a vehicle passes through two intersection points successively, there are two possible scenarios since the length of the vehicle and the distance between the intersection points are unknown, as shown in 4(d). A long truck may cover both intersection points simultaneously, while a short car can be in between.
Example 5 (Overtake on the opposite side)
TSL can use overlapping to create overtaking on opposite scenarios, as shown in 4(e). It contains two roads: one road has two lanes, l1 and l2, running from left to right, and the other road has one lane, l3, running from right to left, with l3 overlapping with l2. Vehicles c1 (blue) and c2 (red) are initially located on l1, and c3 (yellow) is located on l3. With no lane changes allowed for c2 and c3, two scenarios are provided by TSL for c1 to overtake c2: waiting for c3 to pass before overtaking, and overtaking before c3 passes.
Example 6 (The T-intersection)
Figure 2 depicts a typical T-intersection, which combines the primitives shown in the previous examples. TSL reasoning still applies, but there are too many possible scenarios to list here. For the complete example, please refer to our source code.
The examples above show how TSL works for scenario generation. By combining these primitives, TSL can manage different urban road networks and create a variety of traffic scenarios. Please refer to our source code for the detailed implementation of these examples.
The scenarios generated can be described using a scenario description language and simulated as test cases for autonomous vehicles. For instance, Example 1 can serve as a test case for the overtaking maneuver of an autonomous driving system when the system takes control of c1. By examining the system’s behavior in various scenarios, we can assess whether it might fail to respond appropriately to c2. We have demonstrated the effectiveness of this workflow in highway scenarios in our previous study (Wang et al., 2023), and TSL assists in its expansion to general urban scenarios.
In addition to scenario generation, TSL can be used in other ways as an effective reasoning tool for traffic scenarios. We have developed a framework for highway decision-making of autonomous vehicles using a modified version of TSL [^3] This work is under submission to ITSC 2024. Link of the project: https://github.com/Autonomous-Driving-Safety-Project/TSL_Driver. Also, by formally modeling the control of an autonomous vehicle, we can verify the safety of control sequences using TSL, which will be one of our future works.
7 Conclusion
In this paper, we propose a novel spatial-temporal logic, Traffic Scenario Logic, for modeling and reasoning of urban pedestrian-free traffic scenarios. The compatibility with OpenDRIVE, the de facto industry standard of high-definition maps, and the capability to distinguish different homotopy classes of scenarios set TSL apart from existing approaches. These features make TSL well-suited for developing a complete and human-interpretable test scenario library, which is urgently needed by scenario-based autonomous vehicle safety assessment. We provide examples to illustrate TSL’s effectiveness in generating scenarios on various road networks and its capability in generating comprehensive scenarios under complex urban road networks. Our future work will involve incorporating pedestrians into scenarios and exploring additional applications of TSL.